OpenStudy (anonymous):

Differenciate h(x) = x^3 (3x-5)^2 using power and product rule (also chain)

OpenStudy (campbell_st):

use the product rule u = x^3 du/dx = 3x^2 \[v = (3x-5)^2\] \[dv/dx = 2(3x-5) \times 3\] dv/dx = 6(3x - 5) \[dy/dx = du/dx \times v + dv/dx \times u\] dy/dx = 3x^2 (3x-5)^2 + 6(3x-5) x^3 dy/dx = x^2(3x -5)(9x - 15 + 6x) = x^2(3x -5)(15x - 15)

OpenStudy (anonymous):

for 2(3x-5) (3) (x^3) what i did is (6x^3 ) (3x-5) so then when i expand it, the answer is different from urs..but i dont see any mistake doing this way... what did i do wrong?

OpenStudy (anonymous):

and since the previous one has 3x^2 ..so i took out the common factor which is 3x^2

OpenStudy (campbell_st):

well I found x^2(3x-5) as common factor... then look at what was left

OpenStudy (campbell_st):

yep... you can do that..

OpenStudy (anonymous):

but i got 9x^2) ( 3x-5) (x-1) as my answer..

OpenStudy (anonymous):

(3x^2) (3x+5) [(3x-5 + 2)]

OpenStudy (anonymous):

aand then (3x^2) (3x+5) 3x-3 which the answer is (9x^2) (3x+5)(x-1)

OpenStudy (campbell_st):

taking 3x^2(3x - 5) leaves (3x - 5 + 2x) 3x^2(3x -5)(5x - 5) and you could now take 5 as a common factor from the last term 15x^2(5x - 3)(x -1)

OpenStudy (campbell_st):

its 2x and not 2.... x^3 = x^2 and x

OpenStudy (anonymous):

ohhhh!!!! thank you so much!

OpenStudy (anonymous):

campbell how u write the powers as we write in book without using ^